OpenStudy (anonymous):

How many grams of a stock solution that is 92.5 percent H2SO4 by mass would be needed to make 250 grams of a 35.0 percent by mass solution? Show all of the work needed to solve this problem.

OpenStudy (somy):

just by mass word drives me crazy lol let me check

OpenStudy (somy):

here is the formula

OpenStudy (aaronq):

because the concentrations given are the same, you can use the thinking applied to \(C_1V_1=C_2V_2\). So the formula you would use is \(C_1m_1=C_2m_2\) where C is concentration and m is mass.

OpenStudy (somy):

so that means % by mass= (Mr of compound* mole of element in compound) / (Mr of compound) * 100

OpenStudy (sweetburger):

that beautiful cursive though

OpenStudy (anonymous):

So I have to find Mr of H2SO4?

OpenStudy (somy):

just a sec @aaronq where are concentrations?? i kinda lost u

OpenStudy (aaronq):

92.5 percent H2SO4 by mass is a concentration, it's just given in terms of mass and not volume like you usually see.

OpenStudy (somy):

owhhhhhh i didn't know that

OpenStudy (aaronq):

you can do it the way you're proceeding, and i would suggest you do, but the formula i posted is a shortcut.

OpenStudy (somy):

i don't understand my own post lol @aaronq

OpenStudy (anonymous):

I'm so confused lol

OpenStudy (somy):

so if we go with your formula it'll be 92.5* X= 35.0* 250 ?@aaronq

OpenStudy (somy):

im trying to understand so that im able to explain to u @yesy7 lol

OpenStudy (aaronq):

yep

OpenStudy (somy):

okay then your way is piece of cake lol but the formula that i posted a pic of is piece of pain could you kinda help me understand it? @aaronq

OpenStudy (somy):

@yesy7 apparently percentage by mass can be taken as concentration, and since they gave you 2 % by masses and 1 mass in grams and they ask u to find other mass u can use the formula @aaronq gave C1m1= C2m2 im guessing you are using this because solution is same in both cases?

OpenStudy (anonymous):

Oh okay, seems much easier. So it would be set up like how you posted up there ^^?

OpenStudy (somy):

yup :3

OpenStudy (aaronq):

@Somy If you wanted to do it the other way, you would use the formula in the picture to find moles of 250 grams of a 35.0 percent by mass solution. Then using the same formula, find the moles to find the moles of the stock solution (92.5 percent H2SO4) you need, then convert that to mass.

OpenStudy (somy):

this is a thing.....it sounds easy but...im kinda cursed, i don't get easy stuff let me try what you are saying @aaronq

OpenStudy (anonymous):

So it would look like this 92.5 * X=8750 (idk what to do with the x)...

OpenStudy (somy):

oh this part is math

OpenStudy (somy):

92.5X= 8750 x= 8750/ 92.5 x=

OpenStudy (anonymous):

solve for x?

OpenStudy (somy):

yes

OpenStudy (anonymous):

So it's 94.6

OpenStudy (anonymous):

wow that was much easier

OpenStudy (somy):

ikr lol